OpenStudy (anonymous):

alis car uses 12 gallons of each week . find this rate in qaurts per year

OpenStudy (lucaz):

1 week = 12 gallons 1 month = 4 weeks = 4*12 gallons 1 year = 12 months = 12*4*12 gallons